# Pre-trained Sequential Recommendation Framework via Popularity Dynamics

### v2

@@ 5.8.1. Effect of discounting factor $\gamma$
Metric
R@10
N@10
R@10
N@10
R@10
N@10
PrepRec
0.803
0.591
0.472 0.300 0.489 0.264
SasRec
0.815
0.599
0.437
0.290
0.433
0.245
BERT4Rec
0.816 0.602
0.407
0.249
0.433
0.255
@@
5.8.2. Fine-tune Capability
While we focus on zero-shot cross-domain and cross-application transfers, we also investigate PrepRec’s fine-tune capability. To ensure target datasets are smaller than the source, we further process the target datasets such that they are no more than 10% of the source datasets’ total interactions. After further processing, we follow the same experimental setup in § 5.3. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset and report the results in Table 8. We find that PrepRec , after fine-tuning, outperforms the selected baselines on Office and Epinions by up to 12.9%. This indicates that PrepRec is capable of learning from the limited data and can be further fine-tuned to achieve better performance, which further expands the use case of PrepRec .
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
Fine:2 days; Coarse:10 days $*{}^{*}$
0.782 0.573
0.536
0.344 0.795 0.580
Fine:4 days; Coarse:15 days
0.778
0.553
0.537
0.341
0.790
0.574
Fine:7 days; Coarse:30 days
0.760
0.509
0.526
0.334
0.757
0.543
5.8.3. Effect of Different Time Horizons
Dataset Music Office Epinions
Metric
R@10
N@10
R@10
N@10
R@10
N@10
0.782 0.573 0.536 0.344 0.795 0.580
t-2
0.775
0.553
0.530
0.338
0.728
0.516
t-6
0.730
0.497
0.523
0.326
0.682
0.473
t-12
0.705
0.462
0.507
0.313
0.645
0.445
Table 10. Recommendation results for varying prediction time. Where $t-n$ denotes $n$ finer time periods ago. $t-1$ is the default setting, denoted by $*{}^{\prime}*^{\prime}$. The performance drops as $n$ goes larger, indicating the importance of recent trends in sequential recommendation.
5.8.4. Effect of Prediction Time
In PrepRec, we use the popularity trends up to some time as the context to construct universal item representations. To prevent information leakage, we can only use the popularity dynamics computed prior to the time of the interaction, as illustrated in § 4.1.2 and § 4.1.6. Intuitively, we should compute popularity dynamics as close to the actual interaction time as possible. However, experimenting with this can reveal the importance of recent trends in model performance ( Table 10). Thus, denoting $t-n$ to be $n$ fine time periods ago. We round the long-term horizon according to the change in $t$. Per expectation, we find that $t-6$ and $t-12$ drop performance significantly across all datasets while $t-2$ performs relatively close to $t-1$, revealing that for the experimented datasets, capturing the latest popularity changes is crucial.
@@ 5.9. Discussion
PrepRec shows excellent potential as a pre-trained sequential recommendation model, with competitive performance to the best baseline with a fraction of the model size, and is capable of zero-shot transfer across domains and applications. We argue that PrepRec is particularly useful in the following scenarios: 1) initial sequential model when the data in the domain is sparse; 2) backbone for developing more complex sequential recommenders (i.e., prediction interpolation) 3) online recommendation settings. PrepRec does not explicitly model Item IDs, it can also easily generalize to new items/users with no retraining by just updating the popularity dynamics. We leave this as future work.
@@ 6. Conclusion
Inspired by the impact of pre-trained models, we propose PrepRec , a pre-trained sequential recommendation model that is capable of cross-domain and cross-application zero-shot transfer. We assume no access to auxiliary information (e.g., textual data), and we set a baseline for pre-trained sequential recommenders under the most challenging setting. PrepRec learns to represent items based on their popularity dynamics and computes sequence representations via a popularity dynamics-aware transformer. Through extensive experiments, we empirically show that PrepRec can achieve comparable performance to or even outperform state-of-the-art sequential recommenders on regular sequential recommendation tasks. PrepRec also shows excellent generalizability and can zero-shot transfer to a new domain with minimal performance reduction, indicating that popularity dynamics play an important role in sequential recommendation and are crucial for developing generalizable sequential recommenders.

### v3

@@ 5.5.2. Effect of discounting factor $\gamma$
5.5.3. Effect of Different Time Horizons
Dataset Music Office Epinions
Metric R@10 N@10 R@10 N@10 R@10 N@10
Fine:2 days; Coarse:10 days 0.782 0.573 0.536 0.344 0.795 0.580
Fine:4 days; Coarse:15 days 0.778 0.553 0.537 0.341 0.790 0.574
Fine:7 days; Coarse:30 days 0.760 0.509 0.526 0.334 0.757 0.543
5.6. Fine-tune Capability
@@
Metric R@10 N@10 R@10 N@10 R@10 N@10
PrepRec 0.803 0.591 0.472 0.300 0.489 0.264
SasRec 0.815 0.599 0.437 0.290 0.433 0.245
BERT4Rec 0.816 0.602 0.407 0.249 0.433 0.255
@@
We also investigate PrepRec’s fine-tune capability. To ensure target datasets are smaller than the source, we further process the target datasets such that they are no more than 10% of the source datasets’ total interactions. After further processing, we follow the same experimental setup in § 5.2. We fine-tune PrepRec and retrain the baselines from scratch on the target dataset and report the results in Table 9. We find that PrepRec , after fine-tuning, outperforms the selected baselines on Office and Epinions by up to 12.9%, indicating that PrepRec is capable of learning from the limited data and can be further fine-tuned to achieve better performance.
@@ 5.7. Discussion
PrepRec demonstrates the strong ability for zero-shot transfer. We argue that PrepRec is particularly useful in the following scenarios: 1) initial sequential model when the data in the domain is sparse; 2) backbone for developing more complex sequential recommenders (i.e., prediction interpolation) 3) online recommendation settings.
PrepRec captures the popularity shifts in the sequence and is complementary to state-of-the-art sequential recommenders. It is worth noting that item popularity dynamics might not capture everything in users’ preferences, but we believe they are orthogonal components towards capturing user preferences, which could explain why the interpolation results substantially outperform both PrepRec and the selected state-of-the-art baselines (Table 3).
Additionally, time granularity is also crucial for popularity dynamics, and sequence analysis requires careful consideration of the time horizon. Intuitively, when the dataset time precision is less accurate, i.e., weeks or days, we expect the performance to decrease as the sequential information and popularity dynamics become muddled. If the time precision in the training data increases, we can expect more accurate user sequences and more accurate measures of popularity dynamics. In general, time precision will not significantly impact the performance of PrepRec in most scenarios as in practice, online platforms can record precise time data for each user-item interaction. We will include more discussion in the arXiv version of this paper (Wang et al., 2024).
@@ 6. CONCLUSION
In this paper, using the critical insight of popularity dynamics in the user’s sequence, we developed a novel pre-trained sequential recommendation framework, PrepRec, for the zero-shot, cross-domain setting without any auxiliary information. PrepRec learned transferable, universal item representations via popularity dynamics-aware transformers. We empirically showed that PrepRec can achieve excellent zero-shot transfer to a target domain, comparable to state-of-the-art sequential recommenders trained on the target domain. With extensive within-domain experiments, we found performance gains of 11.8% when we interpolated PrepRec ’s results with state-of-the-art sequential recommenders, indicating that PrepRec is learning complementary information. We posit that popularity dynamics are crucial for developing generalizable sequential recommenders.
As part of future work, we plan to investigate: 1) developing more complex sequential recommenders by using PrepRec as a backbone (i.e., prediction interpolation and auxiliary information), and 2) exploring online recommendation settings.
Acknowledgements.
This work was generously supported by the National Science Foundation (NSF) under grant number 2312561. We also would like to thank the anonymous reviewers for their valuable feedback.